OSSC loses sync when receiving 13.75 Khz H Sync

NewHome Forums OSSC, OSSC Pro and DExx-vd isl OSSC – Discussion and support OSSC loses sync when receiving 13.75 Khz H Sync

Viewing 3 posts - 1 through 3 (of 3 total)
  • Author
  • #55793

    Be Kind please this my first post and still trying to understand the best way to sort this out!
    I have 2 early naughties St Nav Units. Both run Analogue RGB with separate HV Sync. They are respectively NTG1 and NTG2.

    The NTG1 System is to have its display upgraded, from ancient Analogue to more modern Digital Display.

    My setup is simple. Analogue RGBHV from the NTG2 unit is sent by a 1.2m cable to the matching analog input of the OSSC. The OSSC HDMI output goes directly back to the new display driver Board (VS-TY2662-V1-99) and thence to the display (ATN065TN14). The OSSC is running Firmware V.90

    All testing has been carried out on the NTG2 unit which produces 15.73Khz H Sync and 59.9hz V sync. The OSSC syncs perfectly, and upscales the original SDTV display by 2 to 768 x 476 progressive. The Configuration is saved in the OSSC

    The NTG1 unit, when using an identical setup, is connected to the OSSC and syncs for a few seconds before losing the connection. The Displayed information during the short OSSC sync period is 13.75 Khz for H Sync and a Vertical sync rate of 49.72Hz.

    An additional problem is that, for entirely unconnected reasons, the NTG1 turns off after 10 or so seconds.

    The OSSC Configuration has been adjusted by me to work better with the NTG2 Sync rates (15.75 Khz for H Sync and 59.9Hz for V Sync) and I suppose my tweaking has affected the ability of the OSSC to sync with the different information from the NTG1 Unit.

    What I need is a likely starting setup for the 13.57Khz H Sync/49.9Hz V Sync rates of the NTG1, so I can at least get a stable display. I can then start working on the improvements.

    Anyone have any ideas, either with detail, or just to revert the OSSC to its Default settings and start from there?
    thanks and regards


    You could try enabling “AV3 interlacefix” from compatibility menu.


    Many thanks for your response.
    However I now think I was a bit hasty in tasking the OSSC input Sync selection process as an issue.
    Having put both H & V Sync streams onto a Tek Scope, I see that after initially (say 2 secs) pure sync pulses, both of an acceptable size and shape, during which the OSSC syncs fine, both H & V pulses become contaminated by, what I can only call ‘phantom’ pulses.
    The scope has no problem triggering on the authentic pulse and the phantoms are barely visible, but seem to be frequent and random in the time domain. The issue could be caused by a number of factors, so more investigation is needed
    I accept that the OSSC looks for clean, properly defined, Sync pulses and has no errant sync pulse lockout capability – that is not its job – so I will need to dig deeper into the problem at the transmission end of the delivery chain
    More later

Viewing 3 posts - 1 through 3 (of 3 total)
  • You must be logged in to reply to this topic.